Need full answer asap please, thank you so much!
Linda bought a desktop computer and a laptop computer. Before finance charges, the laptop cost
$150
more than the desktop. She paid for the computers using two different financing plans. For the desktop the interest rate was
7%
per year, and for the laptop it was
5.5%
per year. The total finance charges for one year were
$227
. How much did each computer cost before finance charges?
DESKTOP: $?
LAPTOP: $?

Answer by Theo(13342)

let x = the cost of the desktop.
let x + 150 = the cost of the laptop.

the finance charge on the desktop was 7%.
the finance charge on the laptop was 5.5%.

the total finance charge for 1 year was 227.

your equation is .07 * x + .055 * (x + 150) = 227

simplify to get .07 * x + .055 * x + .055 * 150 = 227

combine like terms and simplify further to get .125 * x + 8.25 = 227

subtract 8.25 from both sides of the equation to get .125 * x = 218.75

solve for x to get x = 218.75 / .125 = 1750

that's the cost of the desktop.

the cost of the laptop is 150 more than that = 1900

.07 * 1750 + .055 * 1900 = 227.

solution is confirmed to be good.

desktop cost 1750 and laptop cost 1900 before finance charges.